OpenStudy (anonymous):

Simplify the following product of complex numbers. (8 + i) · (-4 – 3i)

sam (.sam.):

-32-24i-4i-3i^2 -32-28i+3 -29-28i

OpenStudy (anonymous):

(8 + i) · (-4 – 3i) =8(-4) + (-3i)(8)+(i)(-4)+(i)(-3i) =-32-24i-4i+3 =-29-28i

OpenStudy (anonymous):

but i^2=-1 so the answer is -29-(-28)=-1

OpenStudy (anonymous):

no, there is only one i in -29-28i (-1)=i^2 -29-28i=-29+28(i^2)(i) but not -29-(-28)=-1

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

thanks. yall make this look to simple been out of school for years im more of a statistics guy.